OpenStudy (anonymous):

plzz plzz plzz somebody help me!!!!! JOAN SELLS KOAN COFFEE FOR 6.20 PER POUND AND AMARETTO COFFEE THAT SELLS FOR 5.80 A POUND. SHE FINDS THAT BY MIXING THESE TWO BLENDS SHE CREATES A COFFEE THAT SELLS WELL. IF SHE USES 18 LBS OF AMARETTTO COFFEE IN THE BLEND AND WISHES TO SELL THE MIXTURE AT 6.10 PER POUND, HOW MANY LBS OF THE KONA COFFEE SHOULD SHE MIX WITH THE AMARETTO COFFEE?

OpenStudy (anonymous):

6.1*(x+18) = 5.8*18 + 6.2*x

OpenStudy (anonymous):

can u please explain what the variables represent :)

OpenStudy (anonymous):

x is pounds of Kona

OpenStudy (phoenixfire):

6.2 per pound of Koan 5.8 per pound of Amaretto She wants to mix the two and sell for 6.1. Let a denote the pounds of Amaretto and k the pounds of Koan that are used: \[6.1(a+k)=5.8a+6.2k\]She uses 18 lbs of Amaretto (a=18) \[6.1(18+k)=5.8(18)+6.2k\]Now you solve for k to find out how many lbs of koan coffee she needs to use to match the money.

OpenStudy (anonymous):

thank you